If you want to have $427,644 in 7 years, how much money should you put in a savings account today? Assume that the savings acThanks so much everyone! I am at a loss...
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

427,644=P*(1.064)^7

P=427,644/(1.064)^7

=427,644*0.647751764

=$277007.16(Approx).